# algorithmic

### v1

@@ Abstract
1 Introduction
Graph Neural Networks (GNNs) have demonstrated powerful expressiveness in graph representation learning [80, 27]. However, what structural information can be learned via GNN remains an open question [67, 9, 15, 5, 42, 79]. Particularly, scant attention has been directed towards this question in terms of structural information specific to two nodes. A prominent task relying on such the information is link prediction. Although a number of GNN-based link prediction models have been introduced [74, 56, 78, 18, 62, 41, 36, 23], many of them lack thorough investigations into whether their models effectively learn pair-specific structural information.
For example, SEAL [74] and its successors [37, 60, 70, 59] are a family of link prediction methods that attempt to use GNNs to learn pair-specific structural information represented by traditional link heuristics such as Common Neighbors, Katz index [26], etc. SEAL [74] has proven that most link heuristics between two nodes in a graph can be computed approximately within an enclosing subgraph specifically constructed for that two nodes. In essence, SEAL-type methods perform link prediction by using GNNs to classify such enclosing subgraphs, with an expectation that GNNs could learn the structural information equivalent to link heuristics. However, a critical evaluation of this expectation is lacking in the literature. We demonstrate that this expectation does not hold completely.
@@ 1 Introduction
First, the majority of GNNs follow a neighborhood aggregation scheme, where each node’s representation is recursively updated by aggregating the representations of that node and its neighbors [12, 36, 27]. The learned representations are node-wise. It has been recognized that every node’s representation can hardly capture information related to the number of its neighbors. This is due to the nature of the set-based pooling of the aggregation scheme, which inherently ignores the size of the neighborhood set of each node [67, 76].
@@
A general strategy for applying node-wise representations learned by GNNs to downstream multiple-node tasks (e.g., link prediction, graph classification, etc.) is to combine the representations of the nodes involved in these tasks. For link prediction, we find that the combination of two nodes’ representations essentially lacks the ability to capture information related to the number of common neighbors. This is mainly because node-wise representations learned by GNNs inherently lack information about the number of neighbors of each node, and most operations of combining two nodes’ representations (e.g., concatenation, Hadamard production, etc.) also do not contain any behaviors of counting how many common neighbors between two nodes. To empirically verify the above, we examine the link prediction performance of an approach that incorporates traditional link heuristics (e.g., Common Neighbors) into the GNN. The approach yields results either superior or comparable to those obtained by using only GNNs, experimentally supporting our analysis.
@@
Moreover, in our experiments, we find that trainable node embeddings (different from pre-trained node embeddings, we refer to trainable node embeddings as those embeddings that can be optimized during the model training) can enhance the performance of GNN-based link prediction models, and the denser the graph, the stronger the enhancement. In particular, by only utilizing node embeddings in GCN [29] or GAT [61], we are able to surpass many link prediction specific methods on two dense graphs, i.e., ogbl-ddi and ogbl-ppa [21]. Our explanation is as follows. Compared to the model weights of a GNN that are shared across all nodes [29, 12, 27], each trainable node embedding is unique to its respective node. This characteristic of node embeddings can benefit the model. When the training is supervised by positive and negative link samples (i.e., two nodes are not linked), the link state of two nodes in every link sample could be encoded into the node embeddings of that two nodes and their neighboring nodes by the neighborhood aggregation algorithm of the GNN. This would enable each node embedding to remember the relationships of that node to other nodes, allowing the model to know better which two nodes are more likely to be linked or not. Moreover, in the neighborhood aggregation of the GNN, the denser graphs would allow each node to see more other nodes, leading to better learning of node embeddings for link prediction.
@@
The insights gained in this study can help identify and interpret the limitations of existing link prediction methods, potentially directing the search for more robust algorithms. To demonstrate this, we present two case studies: first, we show that SEAL-type methods [74, 60, 70, 59] could not effectively learn information about the number of common neighbors. Second, we show that NBFNet [82] lacks the algorithmic capability to train powerful node embeddings for link prediction. Additionally, we compare the empirical performance of various link prediction methods on OGB datasets. The results can be explained with our insights.
2 Notations and Problem Definition
@@
Without loss of generality, we demonstrate our work on homogeneous graphs. Let $\mathcal{G}=(\mathbb{V},\mathbb{E},\mathbf{X})$ denote a graph $\mathcal{G}$ with $N$ nodes, where $\mathbb{V}$ is the set of nodes, $|\mathbb{V}|=N$, $\mathbb{E}$ is the set of edges, and $\mathbf{X}\in\mathbb{R}^{N\times f}$ is the feature matrix of nodes. The $i$-th row of $\mathbf{X}$ (i.e., $\mathbf{x}_{i}\in\mathbb{R}^{f}$) is the feature vector of node $i$. The adjacency matrix is $\mathbf{A}\in\mathbb{R}^{n\times n}$ in which the $i,j$-th entry (i.e., $a_{i,j}$) is $1$ if an edge exists from node $i$ to $j$ and 00 otherwise. The degree of node $i$ is $\mathrm{deg}(i)=\sum_{j\in\mathbb{V}}a_{i,j}$. The degree of the graph $\mathcal{G}$ is the average degree of all nodes. A set of nodes connected directly to a node $v\in\mathbb{V}$ is the first-order or 1-hop neighborhood set of $v$ and is denoted by $\Gamma_{v}$.
@@
3 Statistical Link Heuristics
In the long history of link prediction research, especially prior to the emergence of neural networks, a variety of statistical link prediction methods have been proposed [40, 58, 32]. These statistical methods, known as heuristic methods or link heuristics, often rely on intuitive rules or empirical observations, and often extract structural information specific to the target node pair. Therefore, we can concrete abstract pair-specific structural information using tangible link heuristics.
@@ 3 Statistical Link Heuristics
3.1 NCN-dependent Link Heuristics
Table 1: NCN-dependent link heuristics between nodes $v,u$.
@@ 3.1 NCN-dependent Link Heuristics
$\mathrm{JA}_{v,u}$ [24] $\frac{|\Gamma_{v}\cap\Gamma_{u}|}{|\Gamma_{v}\cup\Gamma_{u}|}$
$\mathrm{AA}_{v,u}$ [2] $\sum_{z\in\Gamma_{v}\cap\Gamma_{u}}\frac{1}{\mathrm{log}|\Gamma_{z}|}$
$\mathrm{RA}_{v,u}$ [81] $\sum_{z\in\Gamma_{v}\cap\Gamma_{u}}\frac{1}{|\Gamma_{z}|}$
Sorensen index[57] $\frac{2|\Gamma_{v}\cap\Gamma_{u}|}{|\Gamma_{v}|+|\Gamma_{u}|}$
Salton index [51] $\frac{|\Gamma_{v}\cap\Gamma_{u}|}{\sqrt{|\Gamma_{v}|\cdot|\Gamma_{u}|}}$
Hub Promoted index [49] $\frac{|\Gamma_{v}\cap\Gamma_{u}|}{min{(|\Gamma_{v}|,|\Gamma_{u}|})}$
Hub Depressed index [49] $\frac{|\Gamma_{v}\cap\Gamma_{u}|}{max{(|\Gamma_{v}|,|\Gamma_{u}|})}$
Table 1 lists several commonly-used NCN-dependent Link heuristics. As shown, Common Neighbors (CN) is defined as the size of the intersection of the first-order neighborhood sets of two nodes. Jaccard (JA) coefficient [24] normalizes the CN by the size of the union of the two nodes’ neighborhood sets. AdamicAdar (AA) [2] and Resource Allocation (RA) [81] suppress the contribution of nodes by penalizing each node with its degree. Sorensen index [57], Salton index [51], Hub Promoted index [49], and Hub Depressed index [49] incorporate the degree of two target nodes with CN. We can see from Table 1 that these heuristics are highly dependent on the number of common neighbors between two nodes $v,u$ (i.e., $|\Gamma_{v}\cap\Gamma_{u}|$).
@@
In addition to the above, many other link heuristics are NCN-dependent. Cannistraci et al. [7] suggest that the likelihood of two nodes forming a link increases if their common neighbors are members of a strongly inner-linked cohort, termed local-community-links. They introduce a modified version of CN, JA, AA, and RA, denoted as CAR-based. Furthermore, some another link heuristics consider the clustering coefficient of the nodes in counting common neighbors, such as node clustering coefficient ($\sum_{z\in\Gamma_{v}\cap\Gamma_{u}|}C(z)$) [66] and node-link clustering coefficient ($\sum_{z\in\Gamma_{v}\cap\Gamma_{u}|}\frac{|\Gamma_{v}\cap\Gamma_{z}|}{|\Gamma_{z}|-1}\times C(z)+\frac{|\Gamma_{u}\cap\Gamma_{z}|}{|\Gamma_{z}|-1}\times C(z)$) [65], where $C(z)=\frac{2|\mathcal{E}_{j,k}:j,k\in\Gamma_{z},\mathcal{E}_{j,k}\in\mathbb{E}|}{|\Gamma_{z}|\times(|\Gamma_{z}|-1)}$ is the local clustering coefficient of the node $z$.
@@
An important family of link heuristics is those counting all paths between two target nodes, such as Katz index [26], Leicht Holme Newman index [34], and Rooted PageRank [74]. These heuristics are indeed NCN-dependent, where first-order and high-order common neighbors are considered. For example, Katz index ($\mathrm{Katz}_{v,u}=\sum_{l=1}^{\infty}\beta^{l}|\{\mathrm{path}_{v,u}^{(l)}\}|$) [26] weighted sums the number of all paths between two target nodes $v,u$, where |$|\{\mathrm{path}_{v,u}^{(l)}\}|$ is the number of all paths between node $v$ and $u$ with the length of $l$, and $\beta$ is a damping factor. For example, if $l=4$, |$|\{\mathrm{path}_{v,u}^{(4)}\}|$ can be computed by |$|\{\mathrm{path}_{v,u}^{(4)}\}|=\sum_{a\in\Gamma_{v},b\in\Gamma_{u}}\mathrm{CN}_{a,b}$, where $\mathrm{CN}_{a,b}$ is the number of common neighbors between nodes $a,b$.
3.2 non-NCN-dependent Link Heuristics
@@
According to our categorization strategy, there exist a limited number of link heuristics that are non-NCN-dependent, including Shortest Path Distance (SPD), Preferential Attachment ($\mathrm{PA}_{v,u}=|\Gamma_{v}|\times|\Gamma_{u}|$) [4], and SimRank [25]. Notably, SPD and PA do not extract information about the number of common neighbors. SimRank, detailed in Algorithm 1, recursively refines similarity scores between every two nodes by considering the neighboring nodes of the two nodes, where the number of common neighbors is ignored essentially. Specifically, the similarity score $s_{i,j}^{(m)}$ between node $i,j$ in the $m$-th iteration is obtained by averaging the similarity scores between all neighbors of $i$ and $j$ from the $(m-1)$-th iteration, where the information about how many common neighbors between $i,j$ can hardly be encoded into $s_{i,j}^{(m)}$.
Algorithm 1 SimRank [25]
1:  Input: Graph $\mathcal{G}=(\mathbb{V},\mathbb{E})$ ($|\mathbb{V}|=N$), decay factor $C$ ($10<C<$ 1), iterations $K$
@@
4 Aggregation-based GNNs
Most GNNs follow a neighborhood information aggregation algorithm, where the representation of each node in a graph is iteratively updated by aggregating the representations of its neighbors and its own [12, 27]. Formally, the representation of a node $i$ updated by the $l$-th layer of a GNN is

### v2

# Can GNNs Learn Link Heuristics? A Concise Review and Evaluation of Link Prediction Methods

@@ Abstract
I Introduction
Graph Neural Networks (GNNs) have demonstrated powerful expressiveness in graph representation learning [zhou2022network, khemani2024review]. However, what structural information can be learned via GNN remains an open question [xu2018powerful, chen2020can, geerts2021expressiveness, bouritsas2022improving, liu2023egnn, zhiyao2024opengsl]. Particularly, scant attention has been directed towards this question in terms of structural information specific to two nodes. A prominent task relying on such the information is link prediction. Although a number of GNN-based link prediction models have been introduced [zhang2018link, singh2021edge, zhao2022learning, guo2022linkless, wang2023multi, liu2023link, li2024evaluating, huang2024theory], many of them lack thorough investigations into whether their models effectively learn pair-specific structural information.
For example, SEAL [zhang2018link] and its successors [li2020distance, teru2020inductive, yin2022algorithm, tan2023bring] are a family of link prediction methods that attempt to use GNNs to learn pair-specific structural information represented by traditional link heuristics such as Common Neighbors, Katz index [katz1953new], etc. SEAL [zhang2018link] has proven that most link heuristics between two nodes in a graph can be computed approximately within an enclosing subgraph specifically constructed for that two nodes. In essence, SEAL-type methods perform link prediction by using GNNs to classify such enclosing subgraphs, with an expectation that GNNs could learn the structural information equivalent to link heuristics. However, a critical evaluation of this expectation is lacking in the literature. We demonstrate that this expectation does not hold completely.
@@ I Introduction
First, the majority of GNNs follow a neighborhood aggregation scheme, where each node’s representation is recursively updated by aggregating the representations of that node and its neighbors [corso2020principal, li2024evaluating, khemani2024review]. The learned representations are node-wise. It has been recognized that every node’s representation can hardly capture information related to the number of its neighbors. This is due to the nature of the set-based pooling of the aggregation scheme, which inherently ignores the size of the neighborhood set of each node [xu2018powerful, zhang2021labeling].
@@
A general strategy for applying node-wise representations learned by GNNs to downstream multiple-node tasks (e.g., link prediction, graph classification, etc.) is to combine the representations of the nodes involved in these tasks. For link prediction, we find that the combination of two nodes’ representations essentially lacks the ability to capture information related to the number of common neighbors. This is mainly because node-wise representations learned by GNNs lack information about the number of neighbors of each node, and most operations of combining two nodes’ representations (e.g., concatenation, Hadamard production, etc.) also do not contain any behaviors of counting how many common neighbors between two nodes. To empirically verify the above, we examine the link prediction performance of an approach that incorporates traditional link heuristics (e.g., Common Neighbors) into the GNN. The approach yields results either superior or comparable to those obtained by using only GNNs, experimentally supporting our analysis.
TABLE I: Summary of Link Prediction Information and GNN Capabilities
Forms of Information for Link Prediction Effectiveness in Link Prediction Can GNNs Learn It?
[0.5mm]Pair-specific structural information NCN-dependent Link Heuristics1, such as Common Neighbors, Adamic-Adar, etc. (Section III) Highly effective in sparse graphs. Might lose discriminative power in dense graphs. (Section V, VIII) Limited: Struggles to capture NCN specifics, inherently washing out the explicit count of neighbors due to set-pooling aggregation. (Section IV, V)
non-NCN-dependent Link Heuristics, such as Shortest Path Distance, SimRank, etc. (Section III) Varied effectiveness Partially: Similar learning styles to certain heuristics like SimRank. (Section VII)
Trainable Node Embeddings Enhances performance, especially in dense graphs. Dominant in very dense graphs (Section V VI, VIII) Yes: Powerful in encoding pairwise relationships into node embeddings (Section V VI, VIII)
node and edge attributes Useful but sensitive to data quality (noise and missing values). Yes: Effectively fuse attributes through neighborhood aggregation. (Section V, VIII)
  • 1
    NCN refers to the Number of Common Neighbors between a pair of nodes.
@@
Moreover, in our experiments, we find that trainable node embeddings (different from pre-trained node embeddings, we refer to trainable node embeddings as those embeddings that can be optimized during the model training) can enhance the performance of GNN-based link prediction models, and the denser the graph, the stronger the enhancement. In particular, by only utilizing node embeddings in GCN [kipf2017semi] or GAT [gat2018graph], we are able to surpass many link prediction specific methods on two dense graphs, i.e., ogbl-ddi and ogbl-ppa [hu2020open]. Our explanation is as follows. Compared to the model weights of a GNN that are shared across all nodes [kipf2017semi, corso2020principal, khemani2024review], each trainable node embedding is unique to its respective node. This characteristic of node embeddings can benefit the model. When the training is supervised by positive and negative link samples (i.e., two nodes are not linked), the link state of two nodes in every link sample could be encoded into the node embeddings of that two nodes and their neighboring nodes by the neighborhood aggregation algorithm of the GNN. This would enable each node embedding to remember the relationships of that node to other nodes, allowing the model to know better which two nodes are more likely to be linked or not. Moreover, in the neighborhood aggregation of the GNN, the denser graphs would allow each node to see more other nodes, leading to better learning of node embeddings for link prediction.
@@
The insights gained in this study can help identify and interpret the limitations of existing link prediction methods, potentially directing the search for more robust algorithms. To demonstrate this, we present two case studies: first, we show that SEAL-type methods [zhang2018link, teru2020inductive, yin2022algorithm, tan2023bring] could not effectively learn information about the number of common neighbors. Second, we show that NBFNet [zhu2021neural] lacks the algorithmic capability to train powerful node embeddings for link prediction. Additionally, we compare the empirical performance of various link prediction methods on OGB datasets. The results can be explained with our insights.
In this paper, we begin by reviewing traditional link heuristics, which represent various types of pair-specific structural information, and examine the capacity of GNNs to learn this information. The effectiveness of node embeddings in enhancing link prediction is also explored. Section III through VII delve into these aspects in detail, with key remarks and findings summarized in Table I. Section VIII presents an analysis of the limitations in existing methods, followed by Section IX which discusses the implications of our findings for practical applications. Lastly, Sections X and XI offer a discussion on the limitations of our study and conclude the paper, respectively.
II Notations and Problem Definition
@@
Without loss of generality, we demonstrate our work on homogeneous graphs. Let $\mathcal{G}=(\mathbb{V},\mathbb{E},\mathbf{X})$ denote a graph $\mathcal{G}$ with $N$ nodes, where $\mathbb{V}$ is the set of nodes, $|\mathbb{V}|=N$, $\mathbb{E}$ is the set of edges, and $\mathbf{X}\in\mathbb{R}^{N\times f}$ is the feature matrix of nodes. The $i$-th row of $\mathbf{X}$ (i.e., $\mathbf{x}_{i}\in\mathbb{R}^{f}$) is the feature vector of node $i$. The adjacency matrix is $\mathbf{A}\in\mathbb{R}^{n\times n}$ in which the $i,j$-th entry (i.e., $a_{i,j}$) is $1$ if an edge exists from node $i$ to $j$ and 0 otherwise. The degree of node $i$ is $\mathrm{deg}(i)=\sum_{j\in\mathbb{V}}a_{i,j}$. The degree of the graph $\mathcal{G}$ is the average degree of all nodes. A set of nodes connected directly to a node $v\in\mathbb{V}$ is the first-order or 1-hop neighborhood set of $v$ and is denoted by $\Gamma_{v}$.
@@
III Statistical Link Heuristics
In the long history of link prediction research, especially prior to the emergence of neural networks, a variety of statistical link prediction methods have been proposed [liben2003link, srilatha2016similarity, kumar2020link]. These statistical methods, known as heuristic methods or link heuristics, often rely on intuitive rules or empirical observations, and often extract structural information specific to the target node pair. Therefore, we can concrete abstract pair-specific structural information using tangible link heuristics.
@@ III Statistical Link Heuristics
III-A NCN-dependent Link Heuristics
Table II lists several commonly-used NCN-dependent Link heuristics. As shown, Common Neighbors (CN) is defined as the size of the intersection of the first-order neighborhood sets of two nodes. Jaccard (JA) coefficient [jaccard1901etude] normalizes the CN by the size of the union of the two nodes’ neighborhood sets. AdamicAdar (AA) [adamic2003friends] and Resource Allocation (RA) [zhou2009predicting] suppress the contribution of nodes by penalizing each node with its degree. Sorensen index [sorensen1948method], Salton index [salton1983introduction], Hub Promoted index [ravasz2002hierarchical], and Hub Depressed index [ravasz2002hierarchical] incorporate the degree of two target nodes with CN. We can see from Table II that these heuristics are highly dependent on the number of common neighbors between two nodes $v,u$ (i.e., $|\Gamma_{v}\cap\Gamma_{u}|$).
TABLE II: NCN-dependent link heuristics between nodes $v,u$.
@@ III-A NCN-dependent Link Heuristics
$\mathrm{JA}_{v,u}$ [jaccard1901etude] $\frac{|\Gamma_{v}\cap\Gamma_{u}|}{|\Gamma_{v}\cup\Gamma_{u}|}$
$\mathrm{AA}_{v,u}$ [adamic2003friends] $\sum_{z\in\Gamma_{v}\cap\Gamma_{u}}\frac{1}{\mathrm{log}|\Gamma_{z}|}$
$\mathrm{RA}_{v,u}$ [zhou2009predicting] $\sum_{z\in\Gamma_{v}\cap\Gamma_{u}}\frac{1}{|\Gamma_{z}|}$
Sorensen index[sorensen1948method] $\frac{2|\Gamma_{v}\cap\Gamma_{u}|}{|\Gamma_{v}|+|\Gamma_{u}|}$
Salton index [salton1983introduction] $\frac{|\Gamma_{v}\cap\Gamma_{u}|}{\sqrt{|\Gamma_{v}|\cdot|\Gamma_{u}|}}$
Hub Promoted index [ravasz2002hierarchical] $\frac{|\Gamma_{v}\cap\Gamma_{u}|}{min{(|\Gamma_{v}|,|\Gamma_{u}|})}$
Hub Depressed index [ravasz2002hierarchical] $\frac{|\Gamma_{v}\cap\Gamma_{u}|}{max{(|\Gamma_{v}|,|\Gamma_{u}|})}$
@@
In addition to the above, many other link heuristics are NCN-dependent. Cannistraci et al. [cannistraci2013link] suggest that the likelihood of two nodes forming a link increases if their common neighbors are members of a strongly inner-linked cohort, termed local-community-links. They introduce a modified version of CN, JA, AA, and RA, denoted as CAR-based. Furthermore, some another link heuristics consider the clustering coefficient of the nodes in counting common neighbors, such as node clustering coefficient ($\sum_{z\in\Gamma_{v}\cap\Gamma_{u}|}C(z)$) [wu2016link] and node-link clustering coefficient ($\sum_{z\in\Gamma_{v}\cap\Gamma_{u}|}\frac{|\Gamma_{v}\cap\Gamma_{z}|}{|\Gamma_{z}|-1}\times C(z)+\frac{|\Gamma_{u}\cap\Gamma_{z}|}{|\Gamma_{z}|-1}\times C(z)$) [wu2016predicting], where $C(z)=\frac{2|\mathcal{E}_{j,k}:j,k\in\Gamma_{z},\mathcal{E}_{j,k}\in\mathbb{E}|}{|\Gamma_{z}|\times(|\Gamma_{z}|-1)}$ is the local clustering coefficient of the node $z$.
@@
An important family of link heuristics is those counting all paths between two target nodes, such as Katz index [katz1953new], Leicht Holme Newman index [leicht2006vertex], and Rooted PageRank [zhang2018link]. These heuristics are indeed NCN-dependent, where first-order and high-order common neighbors are considered. For example, Katz index ($\mathrm{Katz}_{v,u}=\sum_{l=1}^{\infty}\beta^{l}|\{\mathrm{path}_{v,u}^{(l)}\}|$) [katz1953new] weighted sums the number of all paths between two target nodes $v,u$, where $|\{\mathrm{path}_{v,u}^{(l)}\}|$ is the number of all paths between node $v$ and $u$ with the length of $l$, and $\beta$ is a damping factor. For example, if $l=4$, $|\{\mathrm{path}_{v,u}^{(4)}\}|$ can be computed by $|\{\mathrm{path}_{v,u}^{(4)}\}|=\sum_{a\in\Gamma_{v},b\in\Gamma_{u}}\mathrm{CN}_{a,b}$, where $\mathrm{CN}_{a,b}$ is the number of common neighbors between nodes $a,b$.
III-B non-NCN-dependent Link Heuristics
@@
According to our categorization strategy, there exist a limited number of link heuristics that are non-NCN-dependent, including Shortest Path Distance (SPD), Preferential Attachment ($\mathrm{PA}_{v,u}=|\Gamma_{v}|\times|\Gamma_{u}|$) [barabasi2002evolution], and SimRank [jeh2002simrank]. Notably, SPD and PA do not extract information about the number of common neighbors. SimRank, detailed in Algorithm 1, recursively refines similarity scores between every two nodes by considering the neighboring nodes of the two nodes, where the number of common neighbors is ignored essentially. Specifically, the similarity score $s_{i,j}^{(m)}$ between node $i,j$ in the $m$-th iteration is obtained by averaging the similarity scores between all neighbors of $i$ and $j$ from the (m1)(m-1)-th iteration, where the information about how many common neighbors between $i,j$ can hardly be encoded into $s_{i,j}^{(m)}$.
Algorithm 1 SimRank [jeh2002simrank]
1:Input: Graph $\mathcal{G}=(\mathbb{V},\mathbb{E})$ ($|\mathbb{V}|=N$), decay factor $C$ ($0<C<1$), iterations $K$
@@
IV Aggregation-based GNNs
Most GNNs follow a neighborhood information aggregation algorithm, where the representation of each node in a graph is iteratively updated by aggregating the representations of its neighbors and its own [corso2020principal, khemani2024review]. Formally, the representation of a node $i$ updated by the $l$-th layer of a GNN is